Carbon dioxide occupies a 2.54 L containers at 273 K at 100 pascals. What will be the volume when the pressure is 150 Pa and 299 K?
1
Expert's answer
2020-04-23T13:06:27-0400

"V1P1\/T1=V2P2\/T2"

2.54*100/273=x*150/299

X=299*254/273*100= 1.85l

Volume=1.85litres